OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2π). cos 2x - cos x = 0

OpenStudy (aaronq):

um i haven't done this in a while, but i could try to find an identity for cos2x that is simpler to work with

OpenStudy (anonymous):

You probably want to use some trig identities.

OpenStudy (anonymous):

Is it cos squared or cos double x?

OpenStudy (anonymous):

cos double x

OpenStudy (anonymous):

Do you know the double angle formula?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

look it up.

OpenStudy (anonymous):

\[ \cos(2x) =\cos^2(x)-\sin^2(x) \]

OpenStudy (anonymous):

how would i apply that to this problem?

OpenStudy (anonymous):

I'm thinking you want to get rid of the double angle because it makes things tougher.